Restart Night #11 - UT 01 Sep 2020

OSA: S. Allanson

ISp: G. Bechetti

Mtn Manager: B. Bingham

SO: J. Power


Late to open for threatening clouds. Opened for a few hours and completed binocular LUCI detector center data collection and Rotator ZP. Several issues with LFBG transform data collection routine. Humidity and clouds moved in just as we resolved our issues preventing us from taking any LFBG transform data, but RFBG transform data was collected. Rain the second half of the night but no leaks in the chamber. We took N375 Imaging, N1.8 G200, and N1.8 G210 flexure data.


Weather was the biggest issue. IRS error when running IDL transform data collection routine. John restarted. Also LFBG focal startion was hard coded to IRTC2. Doug rectified.

Proposed Plan:

  • PFS Alignment Alignment - daytime - into handover if needed - CRITICAL
  • Focus check time permitting (or if weather keeps us closed at sunset)
  • DONEL1 & L2 Rotator Center (0.5 hr) - high
    • M92, M34, or M52 - fields available for all night
  • DONEL1 Rotator ZP for AGW - (1 hr) - critical
    • Long Slit Method - instructions to be provided by J. Power
      • Suggest doing 3 "masks" to verify the RotZP (based on experience with issues from L2 run)
  • AGW Transform for both LUCIs (use the N30 fieldstop, H or FeII filters) - (2 hr) - high/critical
  • LUCI-1 Imaging check (can run LUCI-2 in parallel to repeat) - high
    • Important to test in and out of field to "stress-test" AGW-1 - does it recover from long throws? and multi-point dithers?
    • Sidereal
      • Dither in-field
        • J1601 (sets around 10pm) - dithering in the field -Kband only
        • NGC 7027 (sets 4am) - cycles through all filters, can stop early
      • Dither out-of-field
        • M52 - (all night) - KJH
        • NGC 40 - (all night) - all filters
    • NS
      • Comet 29/P - (after 10pm) - JHK - need to update for tonight's ephemeris
  • Twilight Flats - low
  • Photometric Standards - (1-2hrs) - medium - only if clear - any seeing
  • AFTER LUCI-1 checkout:
  • LBCs
    • Collimate WRS/DOFPIA - high (4 x 0.25 hrs) - to check LBC-Blue
    • Superfocus sequence -PA=0,180 high (1 hr) - to check LBC-Blue
    • Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr) - to check LBC-Blue
    • Test Science OB - guiding - medium (1hr)
      • Sidereal
      • Non-Sidereal
    • Standard Photometric fields - requires clear skies - any seeing
    • Spectrophotometric Standards - requires clear skies - any seeing
    • Twlight Flats
  • MODS
    • Spectrophotometric stars if clear (and if not clear only if taking data)
    • Photometric Standards - (1 hr) - only if clear - any seeing
    • Mixed Mode LBC-R/MODS-2
    • Skyflats MODS2-Blue
    • Revisit Transient - (1hr) - Med/High
      • Same spectroscopy scripts (run 2x extra eposures - total of 4x900sec each channel)
      • If too faint for spectroscopy, take u,g,r,i,z imaging - script TBD

Closed Dome Activities

  • LUCI-Bino Flexure (requires moving telescope: elevation and RFBG/LFBG rotator)
    • DONEN375 Imaging
    • DONEN180 G200
    • DONEN180 G210 (in progress at end of night)
  • LUCI1/2 Wavelength Calibration

Play by Play

Conditions look boderline before sunset, but threatening clouds hagning around the summit, but clear overhead. Humidity ~90%

1:30 LUCI software was in a funky state. LUCI2 MOS was powered off? Dave powered it back on and restarted the LUCI software. Awake and up and running normally. Verifying fieldstop alignment. MODS fibers have just been swapped, waking mods and taking sievesnaps.

2:25 Threatening clouds are overtaking the summit and humidity is spiking. Setting up to take some FlexComp data

2:42 Setup for flexcomp - OptSpec _Sieve in FPU K filter, rot at -89 to start

N375 Imaging Mirror
  1. Set1 El24deg luci1.20200901.0024-77 luci2.20200901.0009-62 ABORTED! Looks like we are going to open
3:19 Looks like the threatening clouds have moved out and conditions are clear enough to open. there are still some thin clouds hanging around. We will start with the detector center check on M92.

3:30 Steve is open and performing pointing and collimation. I'm bringing up the WFS and going to start cooling the camera in teh mean time in preparation for OSCO later.

3:39 L1 & L2 Rotator Center M92
  • M92 Active preset PA0 luci1.20200901.0078-79, luci1.20200901.0063-64 Script has a typo...its was set to track
  • M92 track preset PA0 luci1.20200901.0080, luci1.20200901.0065

  • M92 Active preset PA0 luci1.20200901.0081-82, luci1.20200901.0066-67 Seeing 1.18" on the guider, DIMM is still unavailable.
  • M92 track preset PA0 luci1.20200901.0083, luci1.20200901.0068
  • M92 track preset PA180 luci1.20200901.0084, luci1.20200901.0069
  • M92 track preset PA270 luci1.20200901.0085, luci1.20200901.0070
  • M92 track preset PA90 luci1.20200901.0086, luci1.20200901.0071
4:08 L1 & L2 RotZP with stnO2020 MOS mask

Seeing is super awful and we have clouds. 1.7-2.2" on the guider Picking a brighter guide star.
  • acquisition luci1.20200901.0087-89, luci1.20200901.0072-74
    LUCI1 X shift 1.03, YShift -2.17, Angle -0.021, RADECError 0.0660as/0.071as
    LUCI2 X shift -0.059, YShift -4.08, Angle 0.106, RADECError 0.080as/0.093as
    Confirmation luci1.20200901.0090, luci1.20200901.0075. Looks great. Lost guide star....
  • Adjusting the RIGHTZEROPOINT to 0.1, LEFTZEROPOINT adjusted in PCSInstrument.conf. Restarted PCS and resending preset.
  • acquisition luci1.20200901.0091-93, luci1.20200901.0076-78. Seeing 1.4-2.1" on guider
    LUCI1 X shift 0.94, YShift -2.24, Angle -0.07 , RADECError 0.07/0.03
    LUCI2 X shift -0.02, YShift -3.99 Angle 0.11** , RADECError 0.09 0.06 Angle hasn't changed so I musn;t be changing the right parameter...
    I didn't save teh update correctly. Whoops. Redoing. Adjusted RIGHTZEROPOINT to 177.15 from 177.05
  • acquisition luci1.20200901.0095-97, luci1.20200901.0080-82. Seeing 1.4" on guider
    LUCI1 X shift 1.03, YShift -2.09, Angle 0.07 , RADECError 0.07/0.07
    LUCI2 X shift -0.36, YShift -4.01 Angle 0.07 , RADECError 0.07 0.09
    luci1.20200901.0098, luci1.20200901.0083
5:17 Seeing has improved a smidgen but is variable 1.0-1.4".. Set up for transform data collendtion on StoneO, binocular. Steve is tweaking pointing before we begin. Setup with N375, N30Fieldstop, H.

List: Stone_O180pm10_luci.list

1 180
11 180
15 180
24 180
23 180
28 180
26 180
33 180
32 180
12 180
10 180

1 190
11 190
15 190
24 190
23 190
28 190
26 190
33 190
32 190

luci2.20200901.0085-105 finishing on GS32 at PA190
We are pretty close to the moon... Background will probably be pretty high on some of these fainter stars.

LUCI1: IRS error... John sasy it is not running on obs4, moving to obs3. But rfbg is chugging along happily

LUCI2 failed on GS23 (entry 5), looks like i accidentally miscounted and restarted the transform collect on item 6.

irs not happy on obs4 although luci2 seemed ok. Moved to obs3. Irs still not running. John restarted irs. Seeing has blown up to 1.5-1.7"

LFBG focal station hard coded as IRTC2. Calling doug for help. He adjusted the focal station. Just as we were starting up, the humidity spiked and threatening clouds moved in

6:18 closed for clouds and high humidity. no LUCI1 transform data collected, lots of LUCI2 transform data....

6:56 Fast moving clouds and spiking humidity, although clear overhead. Conditions volitile and unsure we will be closed for long, but just in case we are I'm setting up to take FlexComp data.

Setup for flexcomp - OptSpec _Sieve in FPU K filter, rot at -89 to start

N375 Imaging Mirror
  1. El90deg luci1.20200901.0120-192 luci2.20200901.0126-198
  2. El68deg luci1.20200901.0192-265 luci2.20200901.0199-271
  3. El46deg luci1.20200901.0266-338 luci2.20200901.0272-344
  4. El24deg luci1.20200901.0339-411 luci2.20200901.0345-417
9:25 Raining on the mountain now .Steve is checking the chamber

9:43 Steve has given the all clear to continue. N1.8 G200 @1.17 zJSpec OptSieve in FPU Argon source
  1. El90deg luci1.20200901.0412-484 luci2.20200901.0418-490
  2. El68deg luci1.20200901.0485-557 luci2.20200901.0491-563
  3. El46deg luci1.20200901.0558-630 luci2.20200901.0564-636
  4. El24deg luci1.20200901.0631-703 luci2.20200901.0637-709
12:07 Steve has given the all clear to continue. N1.8 G210 @1.25 J OptSieve in FPU Argon source
  1. El24deg luci1.20200901.0704-776 luci2.20200901.0710-782
  2. El46deg luci1.20200901.0777-849 luci2.20200901.0783-856 <Dave taking over @01:07UT>
  3. El68deg luci1.20200901.0850-922 luci2.20200901.0857-928
  4. El90deg luci1.20200901.0923-995 luci2.20200901.0929-1002 <Called Geno to start fills when we got to zenith>
UT02:35 0996 1000 1003 1007 90 1x5.0s darks

Stages on AGws are on axis!
L1:1001 L2:1007 before homing

as telescope on obs2:
home -u 1 -m 21
home -u 2 -m 21

L1:1002 L2:1008 after homing

Fills done

N375+G210 @1.25 J OptSieve in FPU Argon source
  1. El24deg luci1.20200901.1003-1075 luci2.20200901.1009-1081
  2. El46deg luci1.20200901.1076-1148 luci2.20200901.1082-1154
  3. El68deg luci1.20200901.1149-1175 luci2.20200901.1155-1181 **
**Transfer to commercial power was not smooth (rotators faulted). This set will need to be re-done from the start.
  1. El90deg luci1.20200901.1xxx-1xxx luci2.20200901.1xxx-1xxx
Waiting for all-clear from mountain.

Restarted SW. Temp control modes set to OPERATION, both sides

Rotators powered on, set to -89deg
LUCIs configured N375+G210 @ J 1.25um, Ar lamp on, lines confirmed
Waiting for temps to stabilize
Also waiting for rollover to new UTDATE @ noon

### N375+G210 @1.25 J
__LUCI1__ LUCI2 EL Comments:___________

0001 0073 0001 0073 90 FlexComp_Calib15.xml
0074 0146 0074 0146 68
0147 0219 0147 0219 46
0220 0292 0220 0292 24

### N375+G200 @1.20 J
__LUCI1__ LUCI2 EL Comments:___________

0293 0365 0293 0365 24 FlexComp_Calib15.xml
0366 0438 0366 0438 46
0439 0511 0439 0511 68
0512 0584 0512 0584 90

0585 0589 0585 0589 darks 1x5.0s

Remaining calibrations needed:
  • N30 imaging (I want to do this at higher least 10deg intervals but it will take >3hr)
  • G150+N375 K / G040+N375 J
  • G150+N180 K / G040+N180 J will need "lamps off" data for G150 K data
  • N180 img
-- BarryRothberg - 31 Aug 2020
Topic revision: r11 - 02 Sep 2020, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback